OpenStudy (nesha97):

9. Write a function with the given characteristics. A polynomial with rational coefficients having roots 3, 3, and 3 − i Help Please

OpenStudy (irishboy123):

complex roots ravel in pairs, known as conjugates so if there is a root \(r_{\alpha} = 3 - i\) then there is a root \(r_{\beta} = 3 + i\) and \((x - r_{\alpha})(x-r_{\beta}) = x^2 - 6x + 10\) if you also have repeated root \(x = 3\), you have \((x+3)^2\) So you have: \((x^2 - 6x + 10 ) (x+3)^2 = x^4 - 17 x^2 + 6 x + 90\)

OpenStudy (bimbim2697):

Hey Irish , are you good at Physics

OpenStudy (irishboy123):

nah mate. i only do polynomial factoring what about you?